rxcklesskaisher:

Which expression is equivalent to 21 + 49? a) 7(3 + 49) b) 7(3 + 7) c) 7(14 + 42) d) 7(14 + 49)

rxcklesskaisher:

@dontsaymyname i need help.

AZ:

what is a common factor between 21 and 49? What numbers multiply to 21 and what numbers multiply to 49?

rxcklesskaisher:

idk

AZ:

Here's an example 12 What numbers multiply to 12? 1 * 12 = 12 2 * 6 = 12 3 * 4 = 12 So all the numbers that multiply to 12 (or we could call them factors) would be 1, 2, 3, 4, 6, and 12

AZ:

So what are all the factors, or all the numbers that can multiply to get you 21?

rxcklesskaisher:

3 7

AZ:

Good, there's also 1 and 21 but you got the important ones for now What about for 49?

rxcklesskaisher:

yes

rxcklesskaisher:

wait is it D

AZ:

You almost got the idea. But we're looking at numbers that multiply to get us the original number. We're not adding/subtracting. Have you seen the distributive property before? a(b + c) = ab + ac We would multiply that number inside. If we take that number outside so like if we have ab + ac We would have to divide each number by `a` so we could put it outside like a(b+c)

rxcklesskaisher:

ok so, it must be what two numbers get me to 21

AZ:

Here's another example 55 + 77 Another way to write 55 is 5*11 Another way to write 77 is 7*11 55 + 77 = (5*11) + (7*11) Both numbers have 11 as a common factor so we can pull it out and we get 11 * (5+ 7) Does that make sense?

rxcklesskaisher:

its b then

daboyjaylen:

Good

AZ:

You got it!

rxcklesskaisher:

finally

rxcklesskaisher:

can i let u know when i need help?

AZ:

Sure. You can make a new post and tag me :)

rxcklesskaisher:

ok

SapphireVoid:

the answer is b because 7 + 3 = 10 and 10 times 7 is 70 so B is your answer